An investigator wishes to study a particular environmental factor, namely, the relationship of cigarette-smoking habits in the parents to the presence or absence of asthma in their oldest child living in the household in the 5- to 9-year-old age range (referred to below as their offspring) Suppose that the investigator finds that (1) if both the mother and father are current smokers then the probability of their offspring having asthma is 0.15 (2) if the mother is a current smoker and the father is not then the probability of their offspring having asthma is 0.13. (3) if the father is a current smoker and the mother is not, then the probability of their offspring having asthma is 0.05 (4) if neither parent is a current smoker, then the probability of their offspring having asthma is 0.04.
The probability that the father is a current smoker is 0.5. If the father is a current smoker then the probability that the mother is a current smoker is 0.6; whereas if the father is NOT a current smoker then the probability that the mother is a current smoker is 0.2. What is the probability that the offspring will have asthma?

Answer by ikleyn(52754)

.

You are given a general info that

(1) if both the mother and father are current smokers 
    then the probability of their offspring having asthma is 0.15 


(2) if the mother is a current smoker and the father is not 
    then the probability of their offspring having asthma is 0.13. 


(3) if the father is a current smoker and the mother is not, 
    then the probability of their offspring having asthma is 0.05 


(4) if neither parent is a current smoker, 
    then the probability of their offspring having asthma is 0.04.


Next, you are given an additional portion of information:

    (a) The probability that the father is a current smoker is 0.5. 
    (b) If the father is a current smoker then the probability that the mother is a current smoker is 0.6;
    (c) whereas if the father is NOT a current smoker then the probability that the mother is a current smoker is 0.2.


From this additional info, it is easy to deduce that

    P(F is a smoker, M is a smoker)     = 0.5*0.6 = 0.3;    (a)

    P(F is a smoker, M is a not smoker) = 0.5*0.4 = 0.2;    (b)

    P(F is not a smoker, M is a smoker) = 0.5*0.2 = 0.1.    (c)


From these occurrences (a), (b) and (c), we deduce that

    P(F is not a smoker, M is not a smoker) = the COMPLEMENT to (a), (b), (c) = 1 - 0.3 - 0.2 - 0.1 = 0.5   (d).


Finally, we combine (1) with (a); (2) with (c), (3) with (b) and (4) with (d), and we get

    P(the offspring will have asthma) = 0.15*0.3 + 0.13*0.1 + 0.05*0.2 + 0.04*0.1 = 0.072.


ANSWER. Based on given info, the probability that the offspring will have asthma is 0.072.

Solved.

Quite easy and straightforward, if to think a bit.